OpenStudy (kayders1997):

Evaluate sin(5x)dx from pi/3 to pi/2

OpenStudy (solomonzelman):

\(\color{black}{\displaystyle \int_{\pi/3}^{\pi/2}\sin(5x)dx }\)

OpenStudy (solomonzelman):

This ??

OpenStudy (kayders1997):

Yes

OpenStudy (solomonzelman):

Are you relevantly new to integration?

OpenStudy (kayders1997):

Yes we're just learning about it Calc 1

OpenStudy (solomonzelman):

Then, set \(\color{black}{\displaystyle u=5x }\)

OpenStudy (solomonzelman):

If u=5x, then du/dx=5. Right?

OpenStudy (solomonzelman):

So, multiplying both sides times dx, we get: du = 5dx

OpenStudy (solomonzelman):

then, solving for dx, we get dx=(1/5)du.

OpenStudy (kayders1997):

Yes

OpenStudy (solomonzelman):

\(\color{black}{\displaystyle \int_{\pi/3}^{\pi/2}\sin(5x)dx }\) \(\color{black}{\displaystyle u=5x }\) \(\color{black}{\displaystyle (1/5)du=dx }\) then, change your limits to fit variable \(u\). (Know how to do this?)

OpenStudy (mathmale):

What Solomon is doing here is using the substitution u=5x. From that he gets du=5dx. Make sure you're comfortable with this before progressing further.

OpenStudy (kayders1997):

Plug pi/2 and pi/3 into u= equation?

OpenStudy (solomonzelman):

Yes!

OpenStudy (mathmale):

But that step comes AFTER you've done the indicated integration. Have you done it?

OpenStudy (solomonzelman):

So if \(\color{black}{\displaystyle x=\pi/2 }\) then \(\color{black}{\displaystyle u=5(\pi/2)=5\pi/2 }\) and for \(\color{black}{\displaystyle x=\pi/3 }\) you get \(\color{black}{\displaystyle u=5(\pi/3)=5\pi/3 }\)

OpenStudy (solomonzelman):

So having made a substitution \(\color{black}{\displaystyle u=5x }\) \(\color{black}{\displaystyle (1/5)du=dx }\) and having accounted for new limits of integration, what is your NEW integral?

OpenStudy (kayders1997):

1/5 sin(u)du from 5pi/3 to 5pi/2?

OpenStudy (solomonzelman):

YES, AWESOME!!!

OpenStudy (solomonzelman):

\(\color{black}{\displaystyle \frac{1}{5} \int_{5\pi/3}^{5\pi/2} \sin(u)du }\)

OpenStudy (solomonzelman):

Can you evaluate this integral?

OpenStudy (solomonzelman):

What is the integral of sin(u) ? (What function, when differentiated, has a derivative of sin(u) ??)

OpenStudy (kayders1997):

-cosu

OpenStudy (solomonzelman):

Yes, fabulous!

OpenStudy (solomonzelman):

\(\large \color{black}{\displaystyle \frac{1}{5} \left.\int_{5\pi/3}^{5\pi/2} \sin(u)du=-\cos(u)\right|_{u=5\pi/3}^{u=5\pi/2} }\)

OpenStudy (solomonzelman):

If you want, you may do the following: \(\large \color{black}{\displaystyle \frac{1}{5} \left.\left.\int_{5\pi/3}^{5\pi/2} \sin(u)du=-\cos(u)\right|_{u=5\pi/3}^{u=5\pi/2}=\cos(u)\right|^{u=5\pi/3}_{u=5\pi/2} }\)

OpenStudy (solomonzelman):

in other words, multiply the whole negative 1, and reverse the bounds. (It still remains equivalent if you think about it.)

OpenStudy (solomonzelman):

the whole `thing by` -1.

OpenStudy (kayders1997):

Alright that may be easier

OpenStudy (solomonzelman):

You can evaluate it either way, though, I am pretty sure;)

OpenStudy (solomonzelman):

go ahead, and take your time:)

OpenStudy (mathmale):

Kayders and Solomon: Keep that (1/5). Looks like the (1/5) was inadvertently dropped.

OpenStudy (kayders1997):

Is it 1/10???

OpenStudy (solomonzelman):

True, mathmale.

OpenStudy (solomonzelman):

Kayders, awesomeness!

OpenStudy (kayders1997):

Thank you so much @SolomonZelman :)

OpenStudy (solomonzelman):

No problem, yw!